# SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating

### v2

@@ Abstract
Transformer models, despite their impressive performance, often face practical limitations due to their high computational requirements. At the same time, such models exhibit significant activation sparsity, which can be leveraged to reduce the inference cost by transforming parts of the network into Mixture-of-Experts (MoE) layers. However, despite the crucial role of activation sparsity, its impact on this process remains unexplored. In this paper, we enhance the efficiency of MoE conversion through activation sparsity enforcement. Moreover, motivated by the high variance in the number of activated neurons, we propose a more effective dynamic-$k$ expert selection rule that adjusts the number of executed experts on a per-token basis. Finally, we extend this approach to multi-head attention projections, which results in even further savings. The proposed method, Sparsified Activation Dynamic-$k$ Mixture-of-Experts (SADMoE), outperforms existing approaches on common NLP and vision tasks, allowing us to save up to 60% of inference cost without significantly affecting model performance.
SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating
[…]
@@ 1 Introduction
footnotetext: ${}^{\ast}$ Equal contribution. Listing order is random. For a more detailed list of contributions please refer to Appendix Gfootnotetext: ${}^{\dagger}$ Correspondence to: bartwojc@gmail.com
Figure 1: Comparison of top-$k$ expert selection in standard MoE and dynamic-$k$ selection in SADMoE. With top-$k$, the same amount of compute is spent on each sample. dynamic-$k$ saves compute on easy inputs and can assign more experts to the hard ones to increase the performance.
@@
Transformer has become a predominant model architecture in various domains of deep learning such as machine translation (Vaswani et al., 2017), language modeling (Devlin et al., 2018; Radford et al., 2019), and computer vision (Dosovitskiy et al., 2020; Khan et al., 2022). The widespread effectiveness of Transformer models in various applications is closely related to their ability to scale efficiently with the number of model parameters (Kaplan et al., 2020), prompting researchers to train progressively larger and larger models (Touvron et al., 2023; Jiang et al., 2023). However, the considerable computational demands of these models often restrict their deployment in practical settings with limited resources.
@@
At the same time, Transformer models exhibit considerable activation sparsity in their intermediate representations Li et al. (2022), which suggests that most of their computations are redundant. Conditional computation methods reduce these unnecessary costs by using only a subset of the model parameters for any given input (Han et al., 2021). In particular, Mixture-of-Experts (MoE) layer (Shazeer et al., 2016), which consists of multiple experts that are sparsely executed for any single input, proved itself as an effective way to decouple the number of parameters of the model from its computational cost (Clark et al., 2022).
@@
In this work, we aim to leverage MoE layers to decrease the inference cost in pre-trained dense models. As shown by Zhang et al. (2022), this can be done by converting dense layers into MoE layers. We demonstrate the relationship between the activation sparsity of the starting model and the efficiency of the final MoE model. We show that computational savings are directly related to sparsity levels, and propose to enforce higher activation sparsity levels before conversion, which leads to substantially improved cost-to-performance trade-off.
@@
We further show that Transformer models exhibit significant variance in the number of activated neurons, which indicates that using standard top-$k$ expert selection is suboptimal. Therefore, we propose a dynamic-$k$ selection scheme (see Figure 1) that adjusts the number of activated experts on a per-token basis. To enable dynamic-$k$ selection, we propose a novel router training scheme that allows us to approximate the relative contribution of each expert, and select only a subset of the most relevant experts for inference. We find that this approach enables the model to efficiently allocate compute between easy and hard inputs, resulting in significant computational gains.
We extend this method to any standalone linear layer such as projections in Multi-Head Attention (MHA) layers, which often account for over 30%  (Song et al., 2022) of overall computations in the model. We propose a replacement procedure, in which every dense layer is substituted by a two-layer MLP module trained to imitate the output of the original layer. This conversion allows us to take advantage of activation sparsity in the replacement modules and further increase the computational gains.
@@
Figure 2: Overview of key components of SADMoE: a) We propose to enhance the activation sparsity in the base model. b) We convert FFN layers in the model to MoE layers with routers that predict the expert contribution. c) We introduce dynamic-$k$ routing that selects the experts based on their relative contribution.
Our experiments on various NLP and vision benchmarks demonstrate that SADMoE achieves significant improvement in terms of cost-to-performance trade-off. Our contributions can be summarized as follows:
  • We demonstrate the impact of increased activation sparsity on the performance of dense-to-sparse-MoE conversion methods.
  • We propose a novel routing objective along with the dynamic-$k$ expert selection scheme that enables the model to select the appropriate number of experts for a given sample.
  • We extend the approach of dense-to-sparse-MoE to attention projections layers, which results in additional computational gains.
2 Related Work
Mixture-of-Experts.
Sparse Mixture-of-Experts layer (MoE) consists of a predefined number of experts along with a gating network that routes each input to a subset of experts. MoE layers were introduced as an efficient way to further increase the capacity of deep neural networks applied for NLP tasks, initially in LSTM models (Shazeer et al., 2016), and later in Transformers (Lepikhin et al., 2020). Since then, they have also been applied to computer vision (Riquelme et al., 2021; Daxberger et al., 2023). MoE layers have gained significant popularity primarily due to their excellent scaling properties (Du et al., 2022; Clark et al., 2022). Nonetheless, training such models is challenging, primarily because gating decisions must be discrete to ensure sparse expert selection. Various methods of training were proposed, some of which include reinforcement learning (Bengio et al., 2013), weighting the expert output by the probability to allow computation of the gradient of the router (Shazeer et al., 2016), or using the Sinkhorn algorithm (Clark et al., 2022). Some of those approaches also suffer from the possibility of load imbalance and therefore require auxiliary losses or alternative expert selection methods (Fedus et al., 2022; Zhou et al., 2022). Interestingly, in many cases fixed routing functions perform similarly to trainable routers (Roller et al., 2021), which suggests that current solutions are largely suboptimal. MoE models can also be derived from pre-trained dense models by splitting the model weights into experts and independently training the routers for each layer (Zhang et al., 2022; Zuo et al., 2022), which avoids most of the problems present in end-to-end training.
Activation sparsity in Transformers.
Li et al. (2022) show that ReLU-based Transformer models produce sparse activations in their intermediate representations, an effect that is prevalent across architectures, layers, and modalities. They propose a simple rule for keeping only top-$k$ activations in each MLP layer, which results in a model with comparable performance. Similarly, Mirzadeh et al. (2023) demonstrate that ReLU activation function in LLMs encourages ensuing activation sparsity that can be leveraged to skip redundant computations. Tuli and Jha (2023) take a step further and design a dedicated Transformer architecture accelerator that also exploits activation sparsity, while Liu et al. (2023) proposes to predict activation sparsity structure in LLMs and reduce the model latency by skipping redundant computations. Jaszczur et al. (2021) demonstrate that it is possible to train Transformer models from scratch with a fixed level of activation sparsity and obtain similar performance. Finally, a related line of works focuses on sparsity in the attention distributions instead of intermediate representations (Correia et al., 2019). None of the above-mentioned methods explore induced activation sparsity as a way to increase computational gains, nor do they address variance of the number of sparse activations on a per-token basis.
@@ 3 Method
SADMoE reduces the computational cost of the model by splitting every MLP module into a MoE layer. In this section, we describe all of its components in detail. The conversion procedure starts with activation sparsity enforcement, which we achieve by fine-tuning the pre-trained model with an auxiliary loss (Section 3.1). The weights in each block are then split into experts (Section 3.2), and the routers are trained to predict the norm of the output of each expert (Section 3.3). During inference, each MoE layer dynamically selects the number of experts to execute for any given sample following a simple dynamic-$k$ rule (Section 3.4). Finally, the whole process can be optionally preceded by MHA projections layer replacement, which allows us to apply the same conversion pipeline (Section 3.5) on the replacement modules. A high-level overview of the entire process is presented in Figure 2.
@@ 3.1 Enforcing activation sparsity
The scheme of reducing inference cost by dividing the model into independently activated modules relies on the well-known phenomenon of activation sparsity exhibited by most deep neural networks (Glorot et al., 2011). Taking inspiration from this observation, we expect that enforcing higher levels of activation sparsity may allow for the execution of an even smaller number of experts, resulting in overall computational savings. To this end, we induce activation sparsity by fine-tuning the model with an additional loss term that enforces higher levels of activation sparsity (Georgiadis, 2019). We apply the square Hoyer regularization (Kurtz et al., 2020; Hoyer, 2004) on the activations of the model:
@@
where $\mathcal{L}_{\text{CE}}$ is cross-entropy loss, and $\alpha$ is the hyperparameter that controls the strength of sparsity enforcement.
@@ 3.2 Expert clustering
We split the two-layer MLP modules into experts using the parameter clustering method proposed by Zhang et al. (2022). Assuming the MLP layers are composed of weights $\bm{W}_{1}$, $\bm{W}_{2}$ and corresponding biases $\bm{b}_{1}$, $\bm{b}_{2}$, we treat the weights of each neuron from $\bm{W}_{1}$ as features and feed them into the balanced $k$-means algorithm (Malinen and Fränti, 2014) that groups neurons with similar weights together. Then, we use the resulting cluster indices to split the first linear layer $\bm{W}_{1}$, the first bias vector $\bm{b}_{1}$, and the second linear layer $\bm{W}_{2}$ into $n$ experts of the same size. The second bias $\bm{b}_{2}$ is not affected by this procedure.
@@ 3.3 Expert contribution routing
Note that Zhang et al. (2022) also train each routing network independently, but their method constructs artificial labels for each input, and then subsequently trains the router as a classifier. We discuss the differences in detail in Appendix A.
@@ 3.4 Dynamic-$k$ gating
Figure 3: Distribution of non-zero activations in the FFN layers in GPT-2-base trained with ReLU on OpenWebText, with and without the sparsity enforcement phase. Both models exhibit significant variance, and the mean-to-variance ratio increases in the sparsified model.
Commonly used MoE layers always execute top-$k$ experts for each token, where $k$ is a predefined hyperparameter. This means that, regardless of the difficulty of the input, the model spends the same amount of compute on each batch (Zhou et al., 2022) or token (Shazeer et al., 2016). While this may be appropriate if the model is trained with the same restriction, we demonstrate why it is suboptimal for a model that was converted from a dense model without retraining. In Figure 3, we plot the distribution of non-zero activations in the intermediate layer of each FFN for a ReLU version of GPT-2, both for the standard model and a model trained with sparsity enforcement as described in Section 3.1. Activations in both models exhibit substantial variance, which indicates that the optimal number of executed neurons varies depending on the input token. Moreover, the sparsification phase increases the variance and further motivates per-token dynamic-$k$ expert selection.
@@
Note that as $\tau$ increases, the number of executed experts decreases along with the overall computational cost. We emphasize that after model deployment $\tau$ can be adjusted without the need for retraining.
@@ 3.5 Conversion of standalone dense layers
Figure 4: A large portion of the computational cost of Transformers is spent on linear projections in Multi-Head Attention. We replace each projection with a small MLP module, which allows us to apply our MoE conversion scheme on them. Attention heads are not shown for clarity of presentation.
A significant amount of computing in deep neural networks is often spent on dense layers that are not followed by any activation function. Dense-to-sparse-MoE conversion methods cannot reduce the costs of such layers due to lack of activation sparsity. This determines an upper bound on the possible computational savings. To overcome it, we substitute dense layers with small MLPs that approximately have the same computational cost and number of parameters. Each MLP is trained to imitate the output of the original dense layer given the same input by minimization of the mean squared error.
In our case, for Transformer architectures, we substitute projection matrices along with their biases in every MHA layer, as depicted in Figure 4. This means that the final model has four MoE layers in the MHA layer and one MoE layer in the FFN layer for each Transformer block. Note that we do not modify the computation of the scaled dot-product attention itself and that this scheme can be applied to any standalone dense layer.
Figure 5: FLOPs vs loss for GPT-2 on OpenWebText for MoEfication and SADMoE. Our method outperforms the baseline at every possible computational budget.
@@ 4 Experiments
To validate the impact of our method, we evaluate its performance on language modeling, text classification, and image classification. We obtain the performance versus computational cost characteristic for each method by evaluating the methods with different inference hyperparameters (either $\tau$ described in Section 3.4 for SADMoE or number of experts $k$ for MoEfication). We report the computational cost of each method in FLOPs [^1] We utilize the fvcore library to count model FLOPs: https://github.com/facebookresearch/fvcore, as it is a device-independent metric that has been shown to correlate well with latency (Mirzadeh et al., 2023).
For MoEfication, we follow the procedure described by Zhang et al. (2022) by converting the activation functions of the pre-trained model to ReLU and then fine-tuning the model. In the case of SADMoE, we also replace activation functions with ReLU, except for Section 5.2, where we demonstrate that our method performs well also with GELU. For the sake of a fair comparison, the total training data budget is always kept the same between different methods. See Appendix F for a more detailed description of the setup. The source code for our experiments is available at: https://github.com/bartwojcik/SADMoE.

### v4

# Exploiting Activation Sparsity with Dense to Dynamic-k Mixture-of-Experts Conversion

@@ Abstract
Transformer models can face practical limitations due to their high computational requirements. At the same time, such models exhibit significant activation sparsity, which can be leveraged to reduce the inference cost by converting parts of the network into equivalent Mixture-of-Experts (MoE) layers. Despite the crucial role played by activation sparsity, its impact on this process remains unexplored. We demonstrate that the efficiency of the conversion can be significantly enhanced by a proper regularization of the activation sparsity of the base model. Moreover, motivated by the high variance of the number of activated neurons for different inputs, we introduce a more effective dynamic-$k$ expert selection rule that adjusts the number of executed experts on a per-token basis. To achieve further savings, we extend this approach to multi-head attention projections. Finally, we develop an efficient implementation that translates these computational savings into actual wall-clock speedup. The proposed method, Dense to Dynamic-$k$ Mixture-of-Experts (D2DMoE), outperforms existing approaches on common NLP and vision tasks, reducing inference cost by up to 60% without significantly impacting performance.
@@ 1 Introduction
Transformers have become a predominant model architecture in various domains of deep learning such as machine translation [47], language modeling [6, 31], and computer vision [7, 21]. The widespread effectiveness of Transformer models in various applications is closely related to their ability to scale efficiently with the number of model parameters [20], prompting researchers to train progressively larger and larger models [45, 19]. However, the considerable computational demands of these models often restrict their deployment in practical settings with limited resources.
@@
At the same time, Transformer models exhibit considerable activation sparsity in their intermediate representations [24], which suggests that most of their computations are redundant. Conditional computation methods can reduce these unnecessary costs by using only a subset of the model parameters for any given input [14]. In particular, Mixture-of-Experts (MoE) layers [38], consisting of multiple experts that are sparsely executed for any input token, are an effective way to decouple the number of parameters of the model from its computational cost [3]. As shown by [52], many pre-trained dense Transformer models can be made more efficient by converting their FFN blocks into MoE layers, a process they call MoEfication.
Contributions of this paper: We consider the following research question: what is the optimal way to convert a generic Transformer model into an equivalent sparse variant? We identify a series of weaknesses of the MoEfication process limiting the resulting accuracy-sparsity tradeoff, and propose corresponding mitigations as follows. We call the resulting algorithm Dense to Dynamic-$k$ Mixture-of-Experts (D2DMoE) and outline it in Figure 1.
  1. 1.
@@
    First, we analyze the relationship between the activation sparsity of the starting model and the efficiency of the final MoE model. We show that computational savings are directly related to sparsity levels, and we correspondingly enforce higher activation sparsity levels before conversion through a lightweight fine-tuning process, which leads to a substantially improved cost-to-performance trade-off.
  2. 2.
    We identify the router training scheme in the original MoEfication algorithm as a limitation of the conversion process. We propose to frame the router training as a regression problem instead, hence our routers directly predict the norm of the output of each expert.
  3. 3.
@@
    We show that Transformer models exhibit significant variance of the number of activated neurons, and standard top-$k$ expert selection in the MoE layers is inefficient. We propose an alternative dynamic-$k$ expert selection scheme that adjusts the number of activated experts on a per-token basis. This approach enables the model to efficiently allocate compute between easy and hard inputs, increasing the overall efficiency.
  4. 4.
    We generalize the conversion method to any standalone linear layer including gated MLP variants commonly found in modern LLMs [45, 42] and projections in Multi-Head Attention (MHA) layers (which often account for over 30% of total computations in Transformer models [39]).
    For MHA, we propose a replacement procedure in which every dense layer is substituted by a two-layer MLP module trained to imitate the output of the original layer.
@@
Figure 1: Key components of D2DMoE: (a) We enhance the activation sparsity in the base model. (b) We convert FFN layers in the model to MoE layers with routers that predict the contribution of each expert. (c) We introduce dynamic-$k$ routing that selects the experts for execution based on their predicted contribution.
We evaluate D2DMoE across benchmarks in text classification, image classification, and language modeling, demonstrating significant improvements in cost-performance trade-offs in all cases. D2DMoE is particularly well-suited for contemporary hardware, as evidenced by our efficient GPU implementation, which we contribute alongside our proposed method.
2 Motivation
MoE models have gained a lot of traction over the last years as an effective architecture to decouple the parameter count from the computational cost of the models [56]. In a MoE layer, hard sparsity is usually enforced explicitly by applying a top-$k$ operation on the outputs of a trainable gating layer. However, many recent works [53, 2, 30] have shown that most Transformers, when trained at scale, build intrinsically sparse and modular representations. Zhang et al. [52] proposed to leverage this naturally emerging modularity with MoEfication - a method that converts dense transformer models into MoE models by grouping FFN weights into experts and subsequently learning small routers that determine which experts to activate. Models converted with MoEfication are able to preserve the performance of the original dense models while using only a fraction of their computational cost. However, we believe that the MoEfication procedure is not optimal, and therefore aim to obtain dense-to-sparse conversion schemes that obtain a better cost-performance trade-off.
(a) Impact of sparsity on MoE conversion
(b) Non-zero activations distribution
(c) Top-$k$ vs dynamic-$k$ gating
Figure 2: (a) Cost-accuracy tradeoff for a MoEfied [27] GPT-2 model obtained starting from models with different levels of activation sparsity. Sparsification correlates with the model performance. (b) Distribution of non-zero activations in the FFN layers in GPT-2-base on OpenWebText, with and without the sparsity enforcement phase. Both models exhibit significant variance, and the mean-to-variance ratio increases in the sparsified model. (c) We propose to exploit the variation in activations through a dynamic-$k$ routing procedure that adapts the number of experts allocated to a sample.
Intuitively, a MoE converted from a sparser base model would be able to perform the original function using a smaller number of experts. To validate this hypothesis, we perform MoEfication on different variants of GPT2-base [^1] We provide the experimental details for this analysis in Section 4.3 and Appendix J. with varying activation sparsity levels and show the results in Figure 2(a). As expected, MoEfication performs better with sparser models. We further investigate the per-token mean and the variance of non-zero neurons in the base and sparsified model, and show the results in Figure 2(b). Observe that different layers use a different number of neurons on average. Moreover, the variance of the number of activated neurons is quite high and becomes even more significant in the sparsified model. This means that static top-$k$ gating as used in MoEfication is not optimal for dense-to-MoE converted models, and a more flexible expert assignment rule that would be able to handle the high per-token and per-layer variance could be beneficial to the efficiency of such models, as illustrated at Figure 2(c). Such dynamic-k gating requires routers that reliably predict the contribution of each expert. We observe that routers obtained through MoEfication do not accurately reflect this contribution. Moreover, their router training procedure depends on the strict sparsity of the model guaranteed by the ReLU activation function. Therefore, we design a novel router training scheme that directly predicts the contribution of each expert and generalizes to the broader family of activation functions. We combine the proposed components (sparsity enforcement, expert contribution routing, and dynamic-$k$ gating) into a single method that we call Dense to Dynamic-$k$ Mixture-of-Experts (D2DMoE), which we describe in detail in the next Section.
@@ 3 Method
D2DMoE reduces the computational cost of the model by splitting every MLP module into a MoE layer. In this section, we describe all of its components in detail. A high-level overview of the entire procedure is presented in Figure 1. The conversion process can be optionally preceded by MHA projection layer replacement (Sec. 3.5), which allows us to apply the same transformation pipeline on all replacement modules.
@@ 3.1 Enforcing activation sparsity
We expect that enforcing higher levels of activation sparsity may allow for the execution of an even smaller number of experts, resulting in overall computational savings. To this end, we induce activation sparsity by fine-tuning the model with an additional loss term that induces activation sparsity [11]. We apply the square Hoyer regularization [22, 17] on the activations of the model:
@@
where $\mathcal{L}_{\text{CE}}$ is cross-entropy loss, and $\alpha$ is the hyperparameter that controls the strength of sparsity enforcement. We find that the pre-trained models recover the original performance with only a fraction of the original training budget (eg. 1B tokens for GPT2-base or Gemma-2B, which is less than 1% of the tokens used for pretraining).
@@ 3.2 Expert clustering
We split the two-layer MLP modules into experts using the parameter clustering method proposed by Zhang et al. [52]. Assuming the MLP layers are composed of weights $\bm{W}_{1}$, $\bm{W}_{2}$ and corresponding biases $\bm{b}_{1}$, $\bm{b}_{2}$, we treat the weights of each neuron from $\bm{W}_{1}$ as features and feed them into the balanced $k$-means algorithm [26] that groups neurons with similar weights together. Then, we use the resulting cluster indices to split the first linear layer $\bm{W}_{1}$, the first bias vector $\bm{b}_{1}$, and the second linear layer $\bm{W}_{2}$ into $n$ experts of the same size. The second bias $\bm{b}_{2}$ is not affected by this procedure.
MoEfication process was designed for standard two-layered MLPs [52]. Recent LLMs [45, 42] have shifted towards gated FFNs, where the activation is realized through a Gated Linear Unit (GLU) [37], which contains an additional weight matrix for the gate projections. To adapt the expert clustering procedure described above to gated FFN layers, we cluster the weights of the gating matrix $\bm{W_{g}}$ instead of $\bm{W_{1}}$, and use the obtained indices to divide the weights of the two other layers. We provide more intuition and details on our method for gated FFNs in Appendix G.
@@ 3.3 Expert contribution routing
Note that Zhang et al. [52] also trains each routing network independently, but their method constructs artificial labels for each input, and then subsequently trains the router as a classifier. We discuss the differences in detail in Appendix A.
@@ 3.4 Dynamic-$k$ gating
Commonly used MoE layers always execute top-$k$ experts for each token, where $k$ is a predefined hyperparameter. This means that, regardless of the difficulty of the input, the model spends the same amount of compute on each batch [54] or token [38]. While this may be appropriate if the model is trained with the same restriction, it is suboptimal for a model that was converted from a dense model, as we show in Section 2.
@@
Note that as $\tau$ increases, the number of executed experts and the overall computational cost decrease. We emphasize that after model deployment $\tau$ can be adjusted without retraining.
@@ 3.5 Conversion of standalone dense layers
A significant amount of computing in deep neural networks is often spent on dense layers that are not followed by any activation function. Dense-to-sparse-MoE conversion methods cannot reduce the costs of such layers due to a lack of activation sparsity. This determines an upper bound on the possible computational savings. To overcome it, we substitute dense layers with small MLPs with approximately the same computational cost and number of parameters. Each MLP is trained to imitate the output of the original dense layer given the same input by minimizing the mean squared error between the two (akin to a distillation loss).
Figure 3: Multi-Head Attention projection conversion scheme.
In our case, for Transformer architectures, we substitute projection matrices along with their biases in every MHA layer, as depicted in Figure 3. This means that the final model has four MoE layers in the MHA layer and one MoE layer in the FFN layer (either plain or gated) for each Transformer block. Note that we do not modify the computation of the scaled dot-product attention itself and this scheme can be applied to any standalone dense layer.
@@ 4 Experiments
To analyze the impact of our method, we evaluate its performance on language modeling, text classification, and image classification. We obtain performance versus computational cost characteristics for each method by evaluating the methods with different inference hyperparameters (either $\tau$ described in Section 3.4 for D2DMoE or number of experts $k$ for MoEfication; we mark them on the plots with dot markers). We report the computational cost of each method in FLOPs, as it is a device-independent metric that has been shown to correlate well with latency [27]. In addition, we measure the wall-clock execution time of an efficient implementation of our method.
For MoEfication, we follow the procedure described by Zhang et al. [52] by converting the activation functions of the pre-trained model to ReLU and then fine-tuning the model. In the case of D2DMoE, we also replace activation functions with ReLU, except for Section 5.4, where we demonstrate that our method performs well also with GELU. To provide a fair comparison, the total training data budget is always the same between different methods. See Appendix J for a detailed description of our setup. The source code for our experiments is available at: https://github.com/bartwojcik/D2DMoE.
(a) ViT-B on ImageNet-1k
(b) BERT-base on CARER
(c) GPT-2-base on OpenWebText
(d) Gemma-2B on C4
Figure 4: FLOPs-performance tradeoff comparison of our method and MoEfication [52] on CV and NLP benchmarks. We also include early-exit (ZTW, [49]) and token dropping baselines (A-ViT, [51]) for classification. Our method outperforms these baselines across multiple computational budgets.
4.1 Image classification
Vision Transfomer [7] is one of the most popular architectures in computer vision. Since our method applies to any Transformer model, we evaluate it on the popular ImageNet-1k [35] dataset. We use a pre-trained ViT-B checkpoint as the base model and compare D2DMoE with MoEfication in terms of the computational cost versus accuracy trade-off. For broader comparison, we also evaluate the state-of-the-art early-exit method Zero-time Waste (ZTW) [49], as well as our re-implementation of A-ViT, an efficient token dropping method proposed by Yin et al. [51]. Our results, presented in Figure 4(a), demonstrate the significant gains from applying our method over MoEfication.
4.2 Text classification
We evaluate our method with BERT-base [6] on the CARER dataset [36] that contains text samples categorized into 6 different emotion categories. We compare the accuracy versus FLOPs trade-off for D2DMoE, MoEfication, and ZTW. We show the results in Figure 4(b). The performance of MoEfication gradually deteriorates and completely collapses when the number of executed experts approaches zero. In comparison, D2DMoE maintains the original performance for a wide range of computational budgets, and the performance drop starts at a significantly lower budget. While early-exiting performs well for the lowest budgets, it obtains worse results than D2DMoE at medium budgets and suffers from a gradual performance decline.